OpenStudy (anonymous):

How many years will it take for a $50 investment to grow to $1000 if it is compounded continuously at a rate of 5%? A(t) = P • e^(r)(t)

OpenStudy (agreene):

1000 = 50 e^(0.05t) Divide both sides by 50 20 = e^0.05t natural log each side ln(20) = ln(e^0.05t) ln(20) = 0.05t divide by 0.05 ln(20)/0.05=t simplify 20ln(20) = t evaluate if you want a real number 59.9146 years = t :)